# Learning from Interaction: A Survey

### v1

@@ Abstract
LLM-based agents can autonomously accomplish complex tasks across various domains. However, to further cultivate capabilities such as adaptive behavior and long-term decision-making, training on static datasets built from human-level knowledge is insufficient. These datasets are costly to construct and lack both dynamism and realism. A growing consensus is that agents should instead interact directly with environments and learn from experience through reinforcement learning. We formalize this iterative process as the Generation-Execution-Feedback (GEF) loop, where environments generate tasks to challenge agents, return observations in response to agents’ actions during task execution, and provide evaluative feedback on rollouts for subsequent learning. Under this paradigm, environments function as indispensable producers of experiential data, highlighting the need to scale them toward greater complexity, realism, and interactivity. In this survey, we systematically review representative methods for environment scaling from a pioneering environment-centric perspective and organize them along the stages of the GEF loop, namely task generation, task execution, and feedback. We further analyze benchmarks, implementation strategies, and applications, consolidating fragmented advances and outlining future research directions for agent intelligence. [^1] We provide a GitHub repository with real-time updates on this topic: https://github.com/lukahhcm/Awesome_Scaling_Environments.
Scaling Environments for LLM Agents in the Era of
Learning from Interaction: A Survey
@@
Figure 1: (A) Experience arises from the Generation-Execution-Feedback (GEF) loop, where environments generate tasks, agents execute them, and environments evaluate and filter useful experience for RL training. (B) Overview of environment scaling: a GEF-aligned taxonomy of environment-scaling methods, alongside implementation, applications, and the unique challenge of Generator-Verifier asymmetry.
@@ 1 Introduction
However, as agent capabilities continue to evolve, it is infeasible to attain intelligence beyond the human-level merely by supervised fine-tuning (SFT) pretrained models on static datasets (huang2025r; su2025crossingrewardbridgeexpanding; zhao2025absolutezeroreinforcedselfplay). Such datasets are typically manually annotated or curated under human oversight, which makes them costly and labor-intensive to produce at scale, intrinsically bounded by human-level knowledge, and lacking realism and adaptability. By contrast, reinforcement learning provides a more aligned training paradigm (tao2024surveyselfevolutionlargelanguage; zhang2025landscapeagenticreinforcementlearning), where agents can explore in the environment, accumulate experiences, and finally acquire new knowledge or skills. We formalize this interactive process as the Generation-Execution-Feedback (GEF) loop, illustrated in Figure 1 (A). In each iteration, the environment first generates diverse tasks, then the agent executes them within the environment, producing action-observation trajectories. The environment subsequently evaluates these rollouts and retains useful experience for subsequent training. Repeated iterations progressively refine the policy and expand the agent’s capabilities. Notably, unlike prior work (gao2025survey), we adopt a broad view of the environment: everything external to the current agent, including the state space, the executable action space, the design of feedback for interaction and evaluation, and the activities of users and other agents, is considered part of it. In this setting, the environment is no longer a mere container for agents’ activities; it has become an active producer of experiential data, underscoring the growing need for scaling environments to create a more complex, realistic, and richly interactive world (camel2025scaling).
Recent research has embraced this trend of scaling the environment from different perspectives. For instance, systems like AgentGen (hu2025agentgen), AgentGym (xi2024agentgym), and GEM (liu2025gemgymagenticllms) devise heterogeneous environments to increase the diversity of the generated tasks. R-Zero (huang2025r) proposes a challenger-solver framework that autonomously generates increasingly difficult tasks. RandomWorld (sullivan2025procedural) scales up the interactivity by procedural generation of diverse tools for agents to access. ARE (andrews2025arescalingagentenvironments) develops an event-driven environment that supports asynchronous interactions between the environment and agents, scaling up the environmental dynamics that conform to realistic settings. However, a systematic analysis that connects these research directions remains absent.
@@
Stage 1: Task Generation (§3) Complexity (§3.1) e.g., ToolLLM (qin2023toolllm), BFCL V3 (patil2025bfcl), $\tau$-bench (yao2024tau), $c^{3}$-bench (yu2025c), TaskCraft (shi2025taskcraft), WebDancer (wu2025webdancerautonomousinformationseeking), WebWalker (wu2025webwalkerbenchmarkingllmsweb), WebSailor (li2025websailornavigatingsuperhumanreasoning), WebShaper (tao2025webshaperagenticallydatasynthesizing), etc. Dynamic  (§3.2) e.g., Eurekaverse (liang2024environment), EvoCurr (cheng2025evocurr), WebRL (qiwebrl), AgentGym (xi2024agentgym), AgentGen (hu2025agentgen), R-Zero (huang2025r), EnvGen (zala2024envgen), ARE (andrews2025arescalingagentenvironments), etc. Diversity  (§3.3) e.g., AgentGym (xi2024agentgym), AgentSense (leng2025agentsense), AgentBank (song2024agentbank), etc. Stage 2: Task Execution (§4) Interactivity (§4.1) e.g., RandomWorld (sullivan2025procedural), AppWorld (trivedi2024appworldcontrollableworldapps), WebShop (yao2023webshopscalablerealworldweb), MCP-Universe (luo2025mcp), MCP-Bench (wang2025mcp), MCPToolBench++ (fan2025mcptoolbench++), AgentScaler (fang2025generalagentic), BrowseMaster (pang2025browsemaster), etc. Realism (§4.2) e.g., RestGPT (song2023restgpt), WebWalker (wu2025webwalkerbenchmarkingllmsweb), AutoMCP (mastouri2025makingrestapisagentready), Tongyi DeepResearch (tongyidr), AgentScaler (fang2025generalagentic), FTRL (ye2025feedbackdriven), $\tau$-bench (yao2024tau), $\tau^{2}$-bench (barres2025tau2bench), APIGen-MT (prabhakar2025apigenmt), ToolHop (ye2025toolhop), Genie 3 (parker2025genie3), Oasis (yang2025oasis), AgentSociety (zhang2025parallelized), ARE (andrews2025arescalingagentenvironments), etc. Stage 3: Feedback  (§5) Density  (§5.1) e.g., VerlTool (jiang2025verltool), Agent-RLVR (da2025agent), Toolrl (qian2025toolrl), SRM (ma2025steplevelrewardmodelsrewarding), ThinkPRM (khalifa2025processrewardmodelsthink), Web-Shepherd (chae2025webshepherd), PR-Clip-Delta (gao2024designing), etc. Granularity  (§5.2) e.g., AdaCtrl (huang2025adactrl), Toolrl (qian2025toolrl), R1-Searcher (song2025r1), RaR (gunjal2025rubricsrewardsreinforcementlearning), Rubicon (huang2025reinforcement), RuscaRL (zhou2025breaking), etc. Automation (§5.3) e.g., JudgeLRM (chen2025judgelrmlargereasoningmodels), GenRM-CoT (zhang2025generativeverifiersrewardmodeling), REWARDAGENT (peng2025agenticrewardmodelingintegrating), ARMAP (chen2025scalingautonomousagentsautomatic), RaR (gunjal2025rubricsrewardsreinforcementlearning), Rubicon (huang2025reinforcement), etc. Objectivity (§5.4) e.g., RM-7B (su2025crossingrewardbridgeexpanding), General-Reasoner (ma2025general), Rlpr (yu2025rlpr), Nemotron-Crossthink (akter2025nemotron), Nover (liu2025nover), Writing-Zero (jia2025writingzerobridgegapnonverifiable), Omni-Thinker (li2025omni), ARE (andrews2025arescalingagentenvironments), etc. Robustness (§5.5) e.g., ScoreDiff (lin2024navigating), MONA (farquhar2025monamyopicoptimizationnonmyopic), PAR (fu2025reward), RRM (liu2024rrm), InfoRM (miao2024inform), RRM (liu2024rrm), RM-7B (su2025crossingrewardbridgeexpanding), Rubicon (huang2025reinforcement), Trinity-RFT (pan2025trinityrftgeneralpurposeunifiedframework), Tongyi DeepResearch (tongyidr), etc.
@@
Therefore, we comprehensively investigate current environment scaling methods and propose a unified taxonomy aligned with the stages of the GEF loop, adopting a pioneering environment-centric perspective. In the task generation stage, we categorize scaling methods into complexity scaling, dynamic scaling, and diversity scaling, which together characterize an environment’s ability to generate challenging, adaptive, and diverse tasks continuously. In the task execution stage, we highlight interactivity and realism, since these properties determine the richness and fidelity of the interaction data from which agents learn. In the feedback stage, we categorize the scaling of evaluative signals along density, granularity, automation, objectivity, and robustness. Beyond this taxonomy, we also analyze current evaluation benchmarks, implementation frameworks, applications, and future research directions. Figure 1 (B) shows a high-level overview of environment scaling, and representative works are listed in Figure 2.
@@
The survey is organized as follows. We first introduce the background and conceptual framework in §2 and §A. We then categorize representative environment scaling methods along the three-stage taxonomy: task generation (§3), task execution (§4), and feedback (§5). Next, we discuss evaluation benchmarks in §B, implementation frameworks in §6, and applications in §C. Finally, we outline future research directions (§7).
2 Background
Scaling Laws for LLM Agents
Generator-Verifier Asymmetry Challenge
A fundamental characteristic in many real-world tasks is the inherent Generator-Verifier Asymmetry (wei2025asymmetry), namely the mismatch between the intelligence required for generator, which generates (§3) or executes (§4) tasks, and that required for verifier, which provides feedback (§5). These two kinds of intelligence naturally form two axes critical to next-generation Agentic AI, as illustrated in Figure 1 (B). From this perspective, scaling up environments essentially corresponds to scaling intelligence along the $x$-axis and the $y$-axis. Current progress in RL largely exploits the regime on the easy-to-verify side of this asymmetry. These Easy-to-Verify, Hard-to-Generate Domains include fields such as mathematics and programming (wei2025browsecomp; jimenez2023swe; phan2025humanity). For these domains, generating and solving a continual stream of high-quality, non-trivial tasks is challenging. In contrast, verification is objective and computationally inexpensive (e.g., via unit tests or exact match on mathematical results). This enables weak-to-strong supervision, where a simple verifier can provide accurate feedback to train a much stronger agent for solving hard tasks. On the contrary, the Hard-to-Verify, Easy-to-Generate Domains include areas such as creative writing, policy-making, or healthcare (lin2024wildbench; arora2025healthbench). For these easy-to-propose, open-ended tasks, verification is subjective, requires substantial expert judgment, or unfolds over long horizons, making high-quality feedback scarce and expensive. This bottleneck, corresponding to the upper-left region of coordinate system, poses more difficulty in modeling the environment, and rendering environment scaling more challenging yet offering greater potential for advancing agent capabilities. Notably, the asymmetry also presents an opportunity: if the generator’s stronger intelligence can be systematically leveraged to strengthen the verifier, so that it can supervise an even stronger generator, then such asymmetric property can be exploited to drive agents’ self-evolution (huang2025r; hong2025cooper; lu2025urpo; chen2025self; wang2025socraticzerobootstrappingreasoning).
@@ 3 Stage 1: Task Generation
In the task generation stage, the environment is required to propose challenging tasks that push the agent toward its capability boundary. Scaling at this stage targets three aspects of the task design: increasing difficulty (complexity scaling § 3.1), introducing dynamics (dynamic scaling § 3.2), and expanding diversity (diversity scaling § 3.3). An illustrative example is shown in Figure 3. For clarity, in complexity scaling, we only consider the intrinsic difficulty of a task (i.e., static complexity). We group the temporal evolution of task difficulty (dynamic complexity) together with changes in the environment itself under the dynamic scaling subsection.
@@ 3.1 Complexity Scaling
Static complexity increases a task’s inherent structural intricacy, moving beyond single-step commands to challenges defined by dependencies, logical flows, and hierarchical relationships. A typical example is LLM tool use, from early single-step tasks to multi-turn, multi-step scenarios, where complexity scales up as the number of turns and steps increases (qin2023toolllm; patil2025bfcl; yao2024tau; yu2025c). More sophisticated tasks exhibit hierarchical or compositional structure, decomposing high-level objectives into nested sub-goals and thereby testing compositional generalization, namely an agent’s ability to solve novel problems by recombining known skills (shao2023compositional). TaskCraft (shi2025taskcraft) operationalizes this by expanding tasks both in depth (longer sequences of tool executions) and in width (multiple sub-goals per objective), enhancing hierarchical reasoning. At the highest level, conditional and graph-based tasks involve non-linear structures with branching logic, where planning must adapt dynamically to intermediate outcomes. Recent efforts in information-seeking agents extend linear sequences to complex graph-based information chains (wu2025webdancerautonomousinformationseeking; tao2025webshaperagenticallydatasynthesizing; li2025websailornavigatingsuperhumanreasoning; wu2025webwalkerbenchmarkingllmsweb), and multi-agent settings further amplify this complexity, as optimal plans become contingent on the actions of other agents, producing intrinsically interdependent, graph-structured challenges.
@@ 3.2 Dynamic Scaling
Task Difficulty Dynamics
Scaling the task’s difficulty dynamically helps agents generalize, which makes the targets non-stationary and changes the actions and states, and we can tune this either on a predetermined schedule or based on how the agent performs. A common strategy is performance-driven scheduling, where the task difficulty is regulated by the success rate (SR), as in Eurekaverse (liang2024environment). Other approaches target newly acquired or weaker skills, as in EvoCurr (cheng2025evocurr) and EnvGen (zala2024envgen). AgentGen’s BI-EVAL mechanism (hu2025agentgen) introduces a bidirectional variation, which adjusts complexity upward or downward to match agent capability, in contrast to earlier methods that mostly increased difficulty (xu2025wizardlmempoweringlargepretrained; luo2025wizardmathempoweringmathematicalreasoning). Beyond these, WebRL (qiwebrl) implements self-adjusting curricula across complex web settings, and AgentGym (xi2024agentgym) generalizes performance-adaptive scheduling to diverse benchmarks. R-Zero (huang2025r) formalizes a challenger-solver paradigm in which a challenger proposes near-boundary tasks based on the solver’s uncertainty, and the solver improves by training on filtered task sets, yielding iterative and targeted curricula.
Environmental Dynamics
Environmental dynamics provide more realistic scenarios for agents. For example, the Meta Agents Research Environments (ARE) platform (andrews2025arescalingagentenvironments) pushes this paradigm further by introducing a more realistic and dynamic environment. In most setups, if the agent is not interacting, the environment typically remains frozen. By contrast, ARE allows the environment to change dynamically through random or scheduled events at all times, allowing it to evolve asynchronously and independently of the agent. As shown in ARE’s Mobile and Gaia2 benchmarks (andrews2025arescalingagentenvironments), agents in such dynamic settings need to balance cognitive depth and temporal responsiveness to manage interaction latency and uncertainty as the environment’s state continually changes. This design shifts the focus from static interaction to more continuous and proactive engagement.

### v3

# Experience Collection: A Survey

@@ Abstract
LLM-based agents can autonomously accomplish complex tasks across various domains. However, to further cultivate capabilities such as adaptive behavior and long-term decision-making, training on static datasets built from human-level knowledge is insufficient. These datasets are costly to construct and lack both dynamism and realism. A growing consensus is that agents should instead interact directly with environments and learn from experience through reinforcement learning. We formalize this iterative process as the Generation-Execution-Feedback (GEF) loop, where environments generate tasks to challenge agents, return observations in response to agents’ actions during task execution, and provide evaluative feedback on rollouts for subsequent learning. Under this paradigm, environments function as indispensable producers of experiential data, highlighting the need to scale them toward greater complexity, realism, and interactivity. In this survey, we systematically review representative methods for environment scaling from a pioneering environment-centric perspective and organize them along the stages of the GEF loop, namely task generation, task execution, and feedback. We further analyze implementation frameworks, challenges, and applications, consolidating fragmented advances and outlining future research directions for agent intelligence.  [^1] We provide a GitHub repository with real-time updates on this topic: https://github.com/lukahhcm/Awesome_Scaling_Environments.
Environment Scaling for Interactive Agentic
Experience Collection: A Survey
@@ 1 Introduction
Figure 1: Experience arises from the Generation-Execution-Feedback (GEF) loop, where environments generate tasks, agents execute them, and environments evaluate and filter useful experience for RL training.
@@
Stage 1: Task Generation (§2) Complexity (§2.1) e.g., ToolBench (qin2023toolllm), BFCL V3 (patil2025bfcl), $\tau$-bench (yao2024tau), $c^{3}$-bench (yu2025c), ToolHop (ye2025toolhop), TaskCraft (shi2025taskcraft), WebDancer (wu2025webdancerautonomousinformationseeking), WebShaper (tao2025webshaperagenticallydatasynthesizing), WebSailor (li2025websailornavigatingsuperhumanreasoning), WebExplorer (liu2025webexplorer), WebLeaper (tao2025webleaper), etc. Dynamic  (§2.2) e.g., EvoCurr (cheng2025evocurr), WebRL (qiwebrl), AgentGym (xi2024agentgym), AgentGen (hu2025agentgen), R-Zero (huang2025r), EnvGen (zala2024envgen), ARE (andrews2025arescalingagentenvironments), RLVE-Gym (zeng2025rlve), etc. Diversity  (§2.3) e.g., TaskCraft (shi2025taskcraft), ARE (andrews2025arescalingagentenvironments), AutoEnv (zhang2025autoenv), AgentGen (hu2025agentgen), AgentGym (xi2024agentgym), AgentBank (song2024agentbank), SWE-Gym (pan2025trainingsoftwareengineeringagents), R2E-Gym (jain2025r2egymproceduralenvironmentshybrid), SWE-Smith (yang2025swesmithscalingdatasoftware), RLVE-Gym (zeng2025rlve), etc. Stage 2: Task Execution (§3) Interactivity (§3.1) e.g., RandomWorld (sullivan2025procedural), AppWorld (trivedi2024appworldcontrollableworldapps), BrowseMaster (pang2025browsemaster), MCP-Bench (wang2025mcp), OSWorld-MCP(jia2025osworldmcp), WebShop (yao2023webshopscalablerealworldweb), FTRL (ye2025feedbackdriven), AndroidWorld (rawles2025androidworld), MiroThinker (miromindteam2025mirothinker), AgentScaler (fang2025generalagentic), etc. Realism (§3.2) e.g., Tongyi DR (tongyidr), AgentScaler (fang2025generalagentic), FTRL (ye2025feedbackdriven), WebShop (yao2023webshopscalablerealworldweb), WebArena (zhou2023webarena), OSWorld (xie2024osworld), OSWorld-G (xie2025scalingcomputerusegroundinguser), $\tau^{2}$-bench (barres2025tau2bench), AgentSociety (zhang2025parallelized), ARE (andrews2025arescalingagentenvironments), etc. Stage 3: Feedback  (§4) Density  (§4.1) e.g., VerlTool (jiang2025verltool), Agent-RLVR (da2025agent), Toolrl (qian2025toolrl), SRM (ma2025steplevelrewardmodelsrewarding), ThinkPRM (khalifa2025processrewardmodelsthink), Web-Shepherd (chae2025webshepherd), PR-Clip-Delta (gao2024designing), etc. Granularity  (§4.2) e.g., AdaCtrl (huang2025adactrl), Toolrl (qian2025toolrl), R1-Searcher (song2025r1), RaR (gunjal2025rubricsrewardsreinforcementlearning), Rubicon (huang2025reinforcement), RuscaRL (zhou2025breaking), DR Tulu (shao2025drtulu), etc. Automation (§4.3) e.g., JudgeLRM (chen2025judgelrmlargereasoningmodels), GenRM-CoT (zhang2025generativeverifiersrewardmodeling), REWARDAGENT (peng2025agenticrewardmodelingintegrating), ARMAP (chen2025scalingautonomousagentsautomatic), RaR (gunjal2025rubricsrewardsreinforcementlearning), Rubicon (huang2025reinforcement), etc. Objectivity (§4.4) e.g., RM-7B (su2025crossingrewardbridgeexpanding), General-Reasoner (ma2025general), Rlpr (yu2025rlpr), BRPO (jia2025writingzerobridgegapnonverifiable), Nemotron-Crossthink (akter2025nemotron), Omni-Thinker (li2025omni), ARE (andrews2025arescalingagentenvironments), etc. Robustness (§4.5) e.g., ScoreDiff (lin2024navigating), MONA (farquhar2025monamyopicoptimizationnonmyopic), PAR (fu2025reward), RRM (liu2024rrm), InfoRM (miao2024inform), RRM (liu2024rrm), RM-7B (su2025crossingrewardbridgeexpanding), Rubicon (huang2025reinforcement), Trinity-RFT (pan2025trinityrftgeneralpurposeunifiedframework), Tongyi DR (tongyidr), etc.
@@
However, as agent capabilities continue to evolve, it is infeasible to attain intelligence beyond the human-level merely by supervised fine-tuning (SFT) pretrained models on static datasets (huang2025r; su2025crossingrewardbridgeexpanding; zhao2025absolutezeroreinforcedselfplay). Such datasets are typically manually annotated or curated under human oversight, which makes them costly and labor-intensive to produce at scale, intrinsically bounded by human-level knowledge, and lacking realism and adaptability. By contrast, reinforcement learning provides a more aligned training paradigm (tao2024surveyselfevolutionlargelanguage; zhang2025landscapeagenticreinforcementlearning), where agents can explore in the environment, accumulate experience, and finally acquire new knowledge or skills. We formalize this interactive process as the Generation-Execution-Feedback (GEF) loop, illustrated in Figure 1. In each iteration, the environment first generates diverse tasks, then the agent executes them within the environment, producing action-observation trajectories. The environment subsequently evaluates these rollouts and retains useful experience for subsequent training. Repeated iterations of this loop progressively refine the policy and expand the agent’s capabilities. In this paradigm, the environment is no longer a mere container for agents’ activities but an active producer of experiential data, underscoring the growing need for scaling environments to create a more complex, realistic, and richly interactive world (camel2025scaling) [^2] Unlike prior work (gao2025survey), we adopt a broad view of the environment: everything external to the current agent, including the state space, the executable action space, the design of feedback for interaction and evaluation, and the activities of users and other agents, is considered part of it. Illustrative examples are shown in Figures 3 and 4..
Recent research has embraced this trend of scaling the environment from different perspectives. Systems like AgentGen (hu2025agentgen) and GEM (liu2025gemgymagenticllms) devise heterogeneous environments to increase the diversity of the generated tasks. R-Zero (huang2025r) proposes a challenger-solver framework that autonomously generates increasingly difficult tasks. RandomWorld (sullivan2025procedural) scales up the interactivity by procedural generation of diverse tools for agents to access. ARE (andrews2025arescalingagentenvironments) develops an event-driven environment that supports asynchronous interactions that conform to realistic settings. However, a systematic analysis that connects these research directions remains absent.
@@
Therefore, we comprehensively investigate current environment scaling methods and propose a unified taxonomy aligned with the stages of the GEF loop, adopting a pioneering environment-centric perspective. In the task generation stage, we categorize methods into complexity scaling, dynamic scaling, and diversity scaling, which together characterize an environment’s ability to generate challenging, adaptive, and diverse tasks continuously. In the task execution stage, we highlight interactivity and realism, since these properties determine the richness and fidelity of the interaction data from which agents learn. In the feedback stage, we categorize the scaling of evaluative signals along density, granularity, automation, objectivity, and robustness. Beyond this taxonomy, we also analyze implementation frameworks, applications, challenges, and future frontiers. Representative works are listed in Figure 2.
@@
The survey is organized as follows. We first categorize representative environment scaling methods along the three-stage taxonomy: task generation (§2), task execution (§3), and feedback (§4), highlighting challenges and frontiers in each dimension. We then analyze implementation frameworks in §5 and future directions in §6. More background (§A), and discussions of method comparisons (B), evaluation benchmarks (§C), and applications (§D) can be found in the Appendix.
@@ 2 Stage 1: Task Generation
In the task generation stage, the environment is required to propose challenging tasks that push the agent toward its capability boundary. Scaling at this stage targets three aspects of the task design: increasing difficulty [^3] For clarity, in complexity scaling, we only consider the intrinsic difficulty of a task (i.e., static complexity). We group the temporal evolution of task difficulty (dynamic complexity) together with changes in the environment itself under the dynamic scaling subsection. (complexity scaling § 2.1), introducing dynamics (dynamic scaling § 2.2), and expanding diversity (diversity scaling § 2.3). An illustrative example is shown in Figure 3.
@@ 2.1 Complexity Scaling
Static complexity increases a task’s inherent structural intricacy, moving beyond single-step commands to challenges defined by dependencies, logical flows, and hierarchical relationships. A typical scenario is tool use, from early single-step tasks to multi-turn, multi-step scenarios, where complexity scales up as the number of turns and steps increases (qin2023toolllm; patil2025bfcl; yao2024tau; ye2025toolhop). More sophisticated tasks exhibit hierarchical or compositional structure, where high-level objectives are decomposed into nested sub-goals, requiring agents to solve novel problems by recombining acquired skills (shao2023compositional; yu2025c; shi2025taskcraft; wu2025webwalkerbenchmarkingllmsweb; wu2025webdancerautonomousinformationseeking). Building on this, TaskCraft (shi2025taskcraft) expands tasks’ structure both in depth (longer sequences of tool executions) and in width (multiple sub-goals per objective), thus enhancing hierarchical reasoning. Further escalating this complexity, conditional and graph-based tasks involve non-linear structures with branching logic, where planning must adapt dynamically to intermediate outcomes (tao2025webshaperagenticallydatasynthesizing; li2025websailornavigatingsuperhumanreasoning; tao2025webleaper; liu2025webexplorer). Such designs lead to significant performance gains, as shown in Table 1, highlighting the importance of scaling up tasks’ complexity to foster robust and generalizable reasoning strategies.
[Uncaptioned image] Future Frontiers
The next step in scaling up task complexity involves: 1) broadening the width by expanding from a limited set of tools to more versatile, multimodal tools, and addressing increasingly complex scenarios that go beyond simple tool invocation and web navigation, and 2) increasing the depth. However, most current errors arise due to context limitations, and synchronous RL frameworks are inefficient at supporting the scaling of task depth, which further exacerbates these issues. Future environmental designs should prioritize supporting flexible, independent updates through asynchronous RL frameworks.
@@ 2.2 Dynamic Scaling
Scaling dynamics focuses on creating an environment that changes task complexity based on agent performance, helping agents learn and adapt to new challenges and goals. This involves scheduling tasks of varying difficulty levels (curriculum learning) based on agent performance metrics, such as success rate (SR) or progress rate (PR), or targeting newly acquired or weaker skills for focused training (zeng2025rlve; liang2024environment; zala2024envgen; qiwebrl; xi2024agentgym; cheng2025evocurr; zala2024envgen). For instance, AgentGen (hu2025agentgen) introduces a bidirectional task variation method, adjusting task complexity upwards or downwards to match the agent’s capabilities, while EnvGen (zala2024envgen) iteratively generates embodied environments that focus on the skills where agents are weak at. More recent advances introduce a challenger-solver co-evolving pipeline, where the challenger proposes near-boundary tasks based on the solver’s uncertainty, and the solver improves by training on filtered task sets, resulting in progressively more difficult curricula and improved reasoning performance (huang2025r). A similar conclusion is drawn by zeng2025rlve, where adaptively shifting a sliding window of difficulty (ensuring appropriately challenging problems throughout training) leads to superior performance gains and improved learning efficiency.
[Uncaptioned image] Future Frontiers
Despite the benefits of adaptively changing environments and task difficulties, most current methods are limited to verifiable tasks or require manual intervention to ensure high-quality task generation (huang2025r; zeng2025rlve). Future designs may focus on developing systematic principles for non-verifiable tasks while also making automated task generation more reliable.